# Language Repository for Long Video Understanding

### v1

@@ Abstract
Language has become a prominent modality in computer vision with the rise of multi-modal LLMs. Despite supporting long context-lengths, their effectiveness in handling long-term information gradually declines with input length. This becomes critical, especially in applications such as long-form video understanding. In this paper, we introduce a Language Repository (LangRepo) for LLMs, that maintains concise and structured information as an interpretable (i.e., all-textual) representation. Our repository is updated iteratively based on multi-scale video chunks. We introduce write and read operations that focus on pruning redundancies in text, and extracting information at various temporal scales. The proposed framework is evaluated on zero-shot visual question-answering benchmarks including EgoSchema, NExT-QA, IntentQA and NExT-GQA, showing state-of-the-art performance at its scale. Our code is available at github.com/kkahatapitiya/LangRepo.
Keywords:
Large Language Models Long-form Video
@@ 1 Introduction
Video data is central to learning systems that can interact and reason about the world. Yet, they also associate with significant challenges such as increased compute requirements and redundant information, to name a few. This is especially critical in long-term videos. Even so, recent literature on video understanding have progressed so far, enabling reasoning capabilities in hours-long video streams [72, 27], in contrast to very-limited temporal spans (e.g. seconds or minutes) just a few years ago. Such progress is intriguing considering how complex the semantics become when temporal span is increased [68, 89]. Work on efficient spatio-temporal attention mechanisms [2, 4], memory management [79, 63], and large-language-models (LLMs) [76, 90, 72] have been key ingredients for such recent improvements.
@@
LLMs, or more-specifically, vision-large-language-models (VLLMs) have been outperforming pure vision models in recent years in all facets, including image-based reasoning [49, 97, 43], grounding [38, 60], video understanding [76, 88, 90], and even robotics [91, 1, 46]. The sheer model scale and the vast pretraining data have enabled such frameworks to capture world knowledge and semantics, beyond what is possible with visual data only. Besides, the ability to process long context-lengths is also key, as it helps modeling long-term dependencies that are crucial for more-complex reasoning and interactions. However, recent studies show that despite the availability of such context-lengths, the effectiveness of models declines with longer input sequences [41]. This promotes the search for alternate representations that can compress input language data without losing meaningful information, essentially managing the context utilization of LLMs.
@@
Moreover, the use of text (i.e., language) in modeling has shown numerous benefits such as rich semantics [77, 52, 31], ease of information sharing between different specialized-models [91] or modalities [49, 21], and interpretability [94, 69]. Among such, interpretability has a huge societal impact in the age of LLMs, to manage adversities such as bias [47, 18] and hallucinations [93, 12]. Simply put, it enables human observers to understand and monitor what really happens within models. Hence, interpretable representations have also been of interest to the community, in place of latent representations [79, 63].
@@
Figure 1: Overview of our Language Repository (LangRepo): We propose an all-textual repository of visual information that updates iteratively, creating a multi-scale and interpretable representation. It extracts information from captions (generated by a VLLM [49, 96]) corresponding to video chunks. In write-to-repo, we prune and rephrase input descriptions, creating concise entries in the repository. In read-from-repo, such language descriptions (together with any optional metadata, e.g., timestamps) at multiple semantic-scales are summarized to generate outputs suited for video VQA. Here, rephrase and summarize are LLM-calls. We also compare LangRepo against state-of-the-art methods, showing strong performance at its scale.
@@
Motivated by the above, we introduce Language Repository (LangRepo), an interpretable representation for LLMs that updates iteratively. It consumes input captions corresponding to video chunks , as shown in Fig. 1 (left). As LangRepo is all-textual, we rely on text-based operations to write and read information. The write operation (write-to-repo) prunes redundant text, creating concise descriptions that keep the context-utilization of LLMs in-check. Its iterative application with increasingly-longer chunks enables it to learn high-level semantics (e.g. long temporal dependencies). The read operation (read-from-repo) extracts such stored language information at various temporal scales, together with other optional metadata within the repository entries (e.g. timestamps). Altogether, our proposed framework is applied to long-term video reasoning tasks including visual question-answering (VQA) on EgoSchema [51], NExT-QA [80] and IntentQA [42], and visually-grounded VQA on NExT-GQA [81], showing strong performance at its scale, as given in Fig. 1 (right). Finally, we ablate our design decisions, providing insights on key components.
@@ 2.0.0.1 Long-video understanding:
Video models have progressed over the years, going from primitive recognition tasks [70, 36] to complex and fine-grained reasoning tasks [68, 89, 80, 22, 51] over long horizons. Both convolutional baselines [6, 17, 16] and transformer architectures [2, 4, 54] have explored research directions such as multi-scale representations [17, 14, 50], efficiency concerns associated with heavy spatio-temporal computations [13, 44], and handling redundant information within video inputs [8, 32]. More recently, long-video understanding has made a leap forward thanks to benchmark datasets [22, 51, 80] and model improvements [90, 92, 56], validating the importance of modeling complex interactions that happen over long periods of time. Still, the sub-par performance of SOTA models on such benchmarks suggests the room for improvement.
@@ 2.0.0.2 Long-context models:
Even before the age of LLMs, models based on convolutions [75, 58, 57, 32], recurrent blocks [24, 10, 26] or transformers [79, 63, 7] have exploited long-term dependencies, especially in the context of video understanding [75, 79] and reinforcement learning [7, 65]. With the rise of LLMs, scaling laws have revealed the importance of longer contexts even more [72, 61], and, thanks to the breakthroughs such as sparse processing [66, 15], caching [20, 37, 33], model-sharding [95, 9, 40], and efficient attention [11, 39], such long-context LLMs have become a reality. Even with very large context lengths, maintaining the effectiveness of reasoning over longer inputs is challenging [41, 85, 67]. This motivates us to think about concise representations that can better-utilize LLM context.
@@ 2.0.0.3 Compressing representations:
When handling heavy inputs, deep learning models have relied on compressed representations. It may come in the form of pruning [64, 5], latent memory [63, 23, 79], or external feature banks [78], to name a few. Despite the intuitive novelties and efficiency gains of such techniques, it is challenging to realize which information get preserved, and how semantically-meaningful they are post-compression. An interpretable representation that supports compression, if available, may shed light on such details.
@@ 2.0.0.4 Language as an interpretable modality:
More-recently, language has emerged as a dominant modality in computer vision due to its strong generalization capabilities [59, 28]. It can also act as a bridge between various domain-specific models [91], other modalities [49, 21], and even human instructions [71, 25], showing intriguing applications in domains such as chat agents (e.g. ChatGPT, Bard) and robot-learning [1, 46]. Since language is interpretable, it enables humans to interact with models more-naturally and make sense of model predictions.
Motivated by the above, we introduce an interpretable language representation that can (1) prune redundant information, and (2) extract multi-scale (or, high-level) semantics, enabling better context-utilization within LLMs.
@@ 3 Observations on long-range inputs
In this section, we investigate how LLMs perform with increasing inputs lengths (i.e., #tokens). Recent LLMs with very-large context lengths such as Gemini-Pro-1.5 [72] (1M tokens) or Claude-2.1 (200k tokens), can support extremely long input sequences. Yet, when feeding longer inputs, the reasoning capabilities (especially, long-term reasoning) of such models diminish. This behavior is also observed in concurrent work [41], and evident in benchmark results of state-of-the-art models [88, 90] (i.e., better performance with shorter inputs, or fewer video frames). To better investigate this in our setup, we evaluate VQA performance on standard long-term video understanding benchmarks while varying the input length (see Table 1). We consider frame/short-clip captions extracted using a VLLM at a baseline framerate (1$\times$) as inputs (introduced in [92]). We either subsample (0.5$\times$) or replicate (2$\times$) the captions, decreasing/increasing the input lengths of a question-answering LLM, namely, Mistral-7B [29] with 8k (or, theoretical 128k) context length. All inputs fit within the context, without any overflow. The observation from this study is consistent: even though the context length of the LLM is sufficient to process given inputs, the effectiveness of its predictions (shown by VQA performance) drops with longer inputs. This motivates us to introduces a concise language representation that preserves important details of long-range inputs, while pruning any redundant information.
Table 1: Observations on increasing input length: We evaluate the performance of a question-answering LLM (Mistral-7B [29]) at different input lengths, on multiple long-video reasoning benchmarks: EgoSchema [51], NExT-QA [80] and IntentQA [42]. Here, 1×1$\times$1 × corresponds to captions extracted by a VLLM at a default frame-rate (and, 0.5×0.5$\times$0.5 ×/2×2$\times$2 × corresponds to a compression/expansion of them by a factor of 2). We see that even though the context length of the LLM can fit all inputs without any overflow, the effectiveness of predictions (i.e., VQA performance) decreases with input length.
Dataset Captions per-video
0.5$\times$
1$\times$
2$\times$
EgoSchema (subset) [51]
49.8
48.8
46.8
NExT-QA [80]
48.2
48.2
46.9
IntentQA [42]
47.1
46.9
45.2
@@ 4 Language Repository
Figure 2: Detailed view of our Language Repository (LangRepo): Here we present the write and read operations within LangRepo. Given short-captions corresponding to video chunks, write-to-repo first prunes redundant captions within each chunk. The same process is iteratively applied on increasingly longer (or, higher-level) chunks— that are already within the repository— to generate multi-scale repository entries. Pruning consists of two stages: (1) grouping most similar captions based on embedding (eg: CLIP [59]) similarities between two subsets, and (2) rephrasing grouped captions with an LLM-call. The resulting LangRepo will include rephrased-captions and any optional metadata (e.g. #occurrences, timestamps). Next, read-from-repo generates concise descriptions for different semantic levels by summarizing the multi-scale language representation, which is also an LLM-call.
We present a Language Repository (LangRepo) that iteratively updates with multi-scale descriptions from video chunks. In contrast to external feature banks [78] or learnable latent memory representations [79, 63, 3], our proposal has a few key advantages: (1) it requires no training (i.e., zero-shot), and (2) it is compatible with both LLM-based processing and human interpretation, as it is fully-textual, i.e., it exists in language-space instead of a latent-space. LangRepo consists of two main operations: (1) information writing (write-to-repo), which prunes redundancies and iteratively updates language descriptions based on increasingly-longer video chunks, and (2) information reading (read-from-repo), which extracts preserved descriptions (with any optional metadata) in multiple temporal scales. We show a detailed view of these operations in Fig. 2, and further elaborate in the following subsections.
Consider a long video that is split in to $n$ non-overlapping chunks, denoted as $V=\{v_{i}\;|\;i=1,\cdots,n\}$. Assume that we already have frame or short-clip captions extracted by a VLLM (e.g. LLaVA [49]) corresponding to such chunks, denoted by $C=\{c_{i}\;|\;i=1,\cdots,n\}$. Here, each chunk may consist of $p$ such captions as in $c_{i}=\{c_{ij}\;|\;j=1,\cdots,p\}$. Altogether, $V$ is represented by $n\times p$ captions which we consider as inputs to our framework.

### v2

@@ Abstract
Language has become a prominent modality in computer vision with the rise of LLMs. Despite supporting long context-lengths, their effectiveness in handling long-term information gradually declines with input length. This becomes critical, especially in applications such as long-form video understanding. In this paper, we introduce a Language Repository (LangRepo) for LLMs, that maintains concise and structured information as an interpretable (i.e., all-textual) representation. Our repository is updated iteratively based on multi-scale video chunks. We introduce write and read operations that focus on pruning redundancies in text, and extracting information at various temporal scales. The proposed framework is evaluated on zero-shot visual question-answering benchmarks including EgoSchema, NExT-QA, IntentQA and NExT-GQA, showing state-of-the-art performance at its scale. Our code is available at github.com/kkahatapitiya/LangRepo.
@@ 1 Introduction
Video data is central to learning systems that can interact and reason about the world. Yet, they also associate with significant challenges such as increased compute requirements and redundant information, to name a few. This is especially critical in long-form videos. Even so, recent literature on video understanding have progressed so far, enabling reasoning capabilities in hours-long video streams (Team et al., 2023; Islam et al., 2024), in contrast to very-limited temporal spans (e.g. seconds or minutes) just a few years ago. Such progress is intriguing considering how complex the semantics become when temporal span is increased (Sigurdsson et al., 2016; Yeung et al., 2018). Work on efficient spatio-temporal attention mechanisms (Arnab et al., 2021; Bertasius et al., 2021), memory management (Wu et al., 2022; Ryoo et al., 2023), and large-language-models (LLMs) (Wang et al., 2022a; Yu et al., 2024; Team et al., 2023) have been key ingredients for such improvements.
@@
LLMs, or more-specifically, vision-large-language-models (VLLMs) have been outperforming pure vision models in recent years in all facets, including image-based reasoning (Liu et al., 2024; Zheng et al., 2024; Li et al., 2023b), grounding (Lai et al., 2023; Rasheed et al., 2023), video understanding (Wang et al., 2022a; Ye et al., 2023b; Yu et al., 2024), and even robotics (Zeng et al., 2022; Ahn et al., 2022; Liang et al., 2023; Li et al., 2024b). The sheer model scale and the vast pretraining data have enabled such frameworks to capture world knowledge and semantics, beyond what is possible with visual data only. Besides, the ability to process long context-lengths is also key, as it helps modeling long-term dependencies that are crucial for more-complex reasoning and interactions. However, recent studies show that despite the availability of such context-lengths, the effectiveness of models declines with longer input sequences (Levy et al., 2024). This promotes the search for alternate representations that can compress input language data without losing meaningful information, essentially managing the context utilization of LLMs.
@@
Moreover, the use of text (i.e., language) in modeling has shown numerous benefits such as rich semantics (Wang et al., 2022b; Menon & Vondrick, 2022; Kahatapitiya et al., 2023), ease of information sharing between different specialized-models (Zeng et al., 2022) or modalities (Liu et al., 2024; Girdhar et al., 2023), and interpretability (Zhao et al., 2023a; Singh et al., 2024). Among such, interpretability has a huge societal impact in the age of LLMs, to manage adversities such as bias (Liang et al., 2021; Ferrara, 2023) and hallucinations (Zhang et al., 2023b; Dhuliawala et al., 2023). Simply put, it enables human observers to understand and monitor what really happens within models. Hence, interpretable representations have also been of interest to the community, in place of latent representations (Wu et al., 2022; Ryoo et al., 2023).
@@
Figure 1: Overview of our Language Repository (LangRepo): We propose an all-textual repository of visual information that updates iteratively, creating a multi-scale and interpretable representation. It extracts information from captions corresponding to video chunks, generated by a VLLM. In write-to-repo, we prune and rephrase input descriptions, creating concise entries in the repository. In read-from-repo, such language descriptions (together with any optional metadata, e.g., timestamps) at multiple semantic-scales are summarized to generate outputs suited for video VQA. Here, rephrase and summarize are LLM-calls. We also compare LangRepo against state-of-the-art methods, showing strong performance at its scale.
@@
Motivated by the above, we introduce Language Repository (LangRepo), an interpretable representation for LLMs that updates iteratively. It consumes input captions corresponding to video chunks , as shown in Fig. 1 (left). As LangRepo is all-textual, we rely on text-based operations to write and read information. The write operation (write-to-repo) prunes redundant text, creating concise descriptions that keep the context-utilization of LLMs in-check. Its iterative application with increasingly-longer chunks enables it to learn high-level semantics (e.g. long temporal dependencies). The read operation (read-from-repo) extracts such stored language information at various temporal scales, together with other optional metadata within the repository entries (e.g. timestamps). Altogether, our proposed framework is applied to long-term video reasoning tasks including visual question-answering (VQA) on EgoSchema (Mangalam et al., 2024), NExT-QA (Xiao et al., 2021) and IntentQA (Li et al., 2023a), and visually-grounded VQA on NExT-GQA (Xiao et al., 2023a), showing strong performance at its scale, as given in Fig. 1 (right). Finally, we ablate our design decisions, providing insights on key components.
@@ 2.0.0.1 Long-video understanding:
Video models have progressed over the years, going from primitive recognition tasks (Soomro et al., 2012; Kuehne et al., 2011) to complex and fine-grained reasoning tasks (Sigurdsson et al., 2016; Yeung et al., 2018; Xiao et al., 2021; Grauman et al., 2022; Mangalam et al., 2024) over long horizons. Both convolutional baselines (Carreira & Zisserman, 2017; Feichtenhofer et al., 2019; Feichtenhofer, 2020) and transformer architectures (Arnab et al., 2021; Bertasius et al., 2021; Nagrani et al., 2021) have explored research directions such as multi-scale representations (Feichtenhofer et al., 2019; Fan et al., 2021; Liu et al., 2022), efficiency concerns associated with heavy spatio-temporal computations (Duke et al., 2021; Li et al., 2019), and handling redundant information within video inputs (Chen et al., 2018; Kahatapitiya & Ryoo, 2021). More recently, long-video understanding has made a leap forward thanks to benchmark datasets (Grauman et al., 2022; Mangalam et al., 2024; Xiao et al., 2021) and model improvements (Yu et al., 2024; Zhang et al., 2023a; Papalampidi et al., 2023), validating the importance of modeling complex interactions that happen over long periods of time. Still, the sub-par performance of SOTA models on such benchmarks suggests the room for improvement.
@@ 2.0.0.2 Long-context models:
Even before the age of LLMs, models based on convolutions (Wang et al., 2018; Piergiovanni & Ryoo, 2018; 2019; Kahatapitiya & Ryoo, 2021), recurrent blocks (Greff et al., 2016; Chung et al., 2014; Hutchins et al., 2022) or transformers (Wu et al., 2022; Ryoo et al., 2023; Chen et al., 2021) have exploited long-term dependencies, especially in the context of video understanding (Wang et al., 2018; Wu et al., 2022) and robotics (Chen et al., 2021; Shang et al., 2022). With the rise of LLMs, scaling laws have revealed the importance of longer contexts even more (Team et al., 2023; Reid et al., 2024), and, thanks to the breakthroughs such as sparse processing (Shazeer et al., 2017; Fedus et al., 2022), caching (Ge et al., 2023; Kwon et al., 2023; Khandelwal et al., 2018), model-sharding (Zhao et al., 2023b; Chowdhery et al., 2023; Lepikhin et al., 2020), and efficient attention (Dao et al., 2022; Lefaudeux et al., 2022), such long-context LLMs have become a reality. Even with very large context lengths, maintaining the effectiveness of reasoning over longer inputs is challenging (Levy et al., 2024; Xiong et al., 2023; Shi et al., 2023). This motivates us to think about concise representations that can better-utilize LLM context.
@@ 2.0.0.3 Compressing representations:
When handling heavy inputs, deep learning models have relied on compressed representations. It may come in the form of pruning (Ryoo et al., 2021; Bolya et al., 2022), latent memory (Ryoo et al., 2023; Graves et al., 2014; Wu et al., 2022), or external feature banks (Wu et al., 2019), to name a few. Despite the intuitive novelties and efficiency gains of such techniques, it is challenging to realize which information gets preserved, and how semantically-meaningful they are post-compression. An interpretable representation that supports compression, if available, may shed light on such details.
@@ 2.0.0.4 Language as an interpretable modality:
More-recently, language has emerged as a dominant modality in computer vision due to its strong generalization capabilities (Radford et al., 2021; Jia et al., 2021). It can also act as a bridge between various domain-specific models (Zeng et al., 2022), other modalities (Liu et al., 2024; Girdhar et al., 2023), and even human instructions (Surís et al., 2023; Gupta & Kembhavi, 2023), showing intriguing applications in domains such as chat agents (e.g. ChatGPT, Bard) and robotics (Ahn et al., 2022; Liang et al., 2023). Since language is interpretable, it enables humans to interact with models naturally and make sense of model predictions.
Motivated by the above, we introduce an interpretable language representation that can (1) prune redundant information, and (2) extract multi-scale (or, high-level) semantics, enabling better context-utilization within LLMs. We rely on open-source LLMs without additional video pretraining, yet showing a strong performance compared to concurrent work based on much-larger proprietary models (Park et al., 2024; Wang et al., 2024b; Fan et al., 2024; Wang et al., 2024e; d; Kim et al., 2024) or video-pertained multi-modal models (Wang et al., 2024a; Li et al., 2024a; Wang et al., 2024c).
@@ 3 Observations on Long-range Inputs
Table 1: Observations on increasing input length: We evaluate the VQA performance of an LLM (Jiang et al., 2023) at different input lengths, on multiple long-video benchmarks (Mangalam et al., 2024; Xiao et al., 2021; Li et al., 2023a). Even with a sufficient context length, the effectiveness of predictions decreases with longer input. Here, 1×1$\times$1 × corresponds to captions generated at a standard frame-rate (and, 0.5×0.5$\times$0.5 ×/2×2$\times$2 × corresponds to a compression/expansion by a factor of 2).
{tabu}
L3.2cmC1.2cmC1.2cmC1.2cm Dataset Captions per-video
0.5$\times$ 1$\times$ 2$\times$
EgoSchema 49.8 48.8 46.8
NExT-QA 48.2 48.2 46.9
IntentQA 47.1 46.9 45.2
In this section, we investigate how LLMs perform with increasing inputs lengths (i.e., #tokens). Recent LLMs with very-large context lengths such as Gemini-Pro-1.5 (Team et al., 2023) (1M tokens) or Claude-2.1 (200k tokens), can support extremely long input sequences. Yet, when feeding longer inputs, the reasoning capabilities (especially, long-term reasoning) of such models diminish. This behavior is also observed in concurrent work (Levy et al., 2024), and evident in benchmark results of state-of-the-art models (Ye et al., 2023b; Yu et al., 2024) (i.e., better performance with shorter inputs, or fewer video frames). To better investigate this in our setup, we evaluate VQA performance on standard long-term video understanding benchmarks while varying the input length (see Table 3). We consider frame/short-clip captions extracted using a VLLM at a baseline framerate (1$\times$) as inputs (introduced in (Zhang et al., 2023a)). We either subsample (0.5$\times$) or replicate (2$\times$) the captions, decreasing/increasing the input lengths of a question-answering LLM, namely, Mistral-7B (Jiang et al., 2023) with 8k (or, theoretical 128k) context length. All inputs fit within the context, without any overflow. The observation from this study is consistent: even though the context length of the LLM is sufficient to process given inputs, the effectiveness of its predictions (shown by VQA performance) drops with longer inputs. This motivates us to introduce a concise language representation that preserves important details of long-range inputs, while pruning any redundant information.
@@ 4 Language Repository
Algorithm Long-video VQA pipeline in LangRepo
require captions of video chunks $C^{0}=\{c_{i}^{0}\;|\;i=1,\cdots,n\}$, number of iterations $K$.
def write-to-repo($c$):
$\quad c_{\text{dst}},\;c_{\text{src}}=\texttt{split}(c)$
$\quad\text{sim}_{\text{src-dst}}=\texttt{similarity}(\texttt{encode}(c_{\text{src}}),\;\texttt{encode}(c_{\text{dst}}))$
$\quad c_{\text{grp}}=\texttt{group}(c_{\text{dst}},\;c_{\text{src}},\;\text{sim}_{\text{src-dst}})$
$\quad{c_{\text{reph}}}=\texttt{rephrase}(\texttt{template}_{\texttt{reph}}(c_{\text{grp}}))$
[…]
return $\quad\textbf{return }r$
def read-from-repo($r$):
$\quad{d}=\texttt{summarize}(\texttt{template}_{\texttt{sum}}(r))$
return $\quad\textbf{return }d$
$r^{0}_{i}=\texttt{write-to-repo}(c_{i}^{0})$
$d^{0}_{i}=\texttt{read-from-repo}(r^{0}_{i})$
for $k\;\text{in range}(K)$:  // iterative write and read
$\quad C^{k+1}=\texttt{re-chunk}([\cdots,\;r^{k}_{i},\;\cdots])$
$\quad r^{k+1}_{i^{\prime}}=\texttt{write-to-repo}(c_{i^{\prime}}^{k+1})$
$\quad d^{k+1}_{i^{\prime}}=\texttt{read-from-repo}(r^{k+1}_{i^{\prime}})$
[…]
Figure 2: Detailed view of our Language Repository (LangRepo): Here we present the write and read operations within LangRepo. Given short-captions corresponding to video chunks, write-to-repo first prunes redundant captions within each chunk. The same process is iteratively applied on increasingly longer (or, higher-level) chunks— that are already within the repository— to generate multi-scale repository entries. Pruning consists of two stages: (1) grouping most similar captions based on embedding (e.g. CLIP (Radford et al., 2021)) similarities between two subsets, and (2) rephrasing grouped captions with an LLM-call. The resulting LangRepo will include rephrased-captions and any optional metadata (e.g. #occurrences, timestamps). Next, read-from-repo generates concise descriptions for different semantic levels by summarizing the multi-scale language representation, which is also an LLM-call.
We present a Language Repository (LangRepo) that iteratively updates with multi-scale descriptions from video chunks. In contrast to external feature banks (Wu et al., 2019) or learnable latent memory representations (Wu et al., 2022; Ryoo et al., 2023; Balažević et al., 2024), our proposal has a few key advantages: (1) it requires no training (i.e., zero-shot), and (2) it is compatible with both LLM-based processing and human interpretation, as it is fully-textual, i.e., it exists in language-space instead of a latent-space. LangRepo consists of two main operations: (1) information writing (write-to-repo), which prunes redundancies and iteratively updates language descriptions based on increasingly-longer video chunks, and (2) information reading (read-from-repo), which extracts preserved descriptions (with any optional metadata) in multiple temporal scales. We show a detailed view of these operations in Fig. 2, and further elaborate in the following subsections.
Consider a long video that is split in to $n$ non-overlapping chunks, denoted as $V=\{v_{i}\;|\;i=1,\cdots,n\}$. Assume that we already have frame or short-clip captions extracted by a VLLM (e.g. LLaVA (Liu et al., 2024)) corresponding to such chunks, denoted by $C^{0}=\{c^{0}_{i}\;|\;i=1,\cdots,n\}$. Here, each chunk may consist of $p$ such captions as in $c^{0}_{i}=\{c^{0}_{ij}\;|\;j=1,\cdots,p\}$. Altogether, $V$ is represented by $n\times p$ captions which we consider as inputs to our framework.